Alice sells her residential investment property for $500,000 that she held for five years. She bought the property for $320,000 and claimed a building depreciation allowance of $12,000 each year. What is her assessable capital gain


1
Expert's answer
2021-11-15T03:12:05-0500

Solution:

Purchase cost of property = Cost - accumulated depriciation

"=320000-(5\times 12000)
\\=320000-60000
\\=260000"

Sale amount consideration = 500000

Then, assessable capital gain = "500000-260000=\$ 240,000"